CASES99: Logbook Entries

CASES99: Hotfilm Messages: 8 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
164 Sat 23-Oct-1999Reset Reference Voltagesmiller
149 Thu 21-Oct-1999Daily Maintenancemiller
140 Tue 19-Oct-1999moved 1.5 sensors to .5mmiller
124 Sun 17-Oct-1999bad film data on 16 OCTmiller
122 Sun 17-Oct-1999hfilms down 1600 1630miller
112 Sat 16-Oct-1999Hotfilm sheath signal is backtowersemmer
111 Sat 16-Oct-1999Signals from Hotfilm sheath down for awhiletowersemmer
85 Tue 12-Oct-1999towermiller


164: Hotfilm, Site , Sat 23-Oct-1999 12:25:22 CDT, Reset Reference Voltages
23 Oct
1200 - 1205 reset bridge reference voltages to 5.00 Volts
1206 performed zero wind calibration for the remainder on the hour using the reset reference voltages.
1207 running


149: Hotfilm, Site , Thu 21-Oct-1999 11:25:19 CDT, Daily Maintenance
21 Oct
1110-1120 CST Reset all reference voltages to 5.00 Volts


140: Hotfilm, Site , Tue 19-Oct-1999 14:18:36 CDT, moved 1.5 sensors to .5m
19 Oct
12:51 tried to blow the cobwebs off the cold film and triple film at 1.5m - partially successful

1300-1400  lowered 1.5 m arm to 0.5m.  Releveled sensors - sonic wouldn't level completely - leaning slightly sideways - toward true north. It appears that the weight of the hf holders on that side has bent the arm slightly.

had to remove the tube from the Spanish pressure sensor to lower the arm.  Resealed the connection with silicone at 13:55.

13:44 Reset all reference voltages on the bridges to 5.00 v. 


124: Hotfilm, Site , Sun 17-Oct-1999 18:11:12 CDT, bad film data on 16 OCT
16 Oct
Checked data for previous day.  Bad film (vu.hf.5m)  on 16 Oct from about 1000 to 2200 local time.



122: Hotfilm, Site , Sun 17-Oct-1999 16:41:37 CDT, hfilms down 1600 1630
17 Oct
1600 - 1630 reset hf/cf bridge voltages and checked levels.  All sensors are leaning slightly to the north (ie level bubble still in central circle).




112: Hotfilm, Site tower, Sat 16-Oct-1999 16:12:28 CDT, Hotfilm sheath signal is back

111: Hotfilm, Site tower, Sat 16-Oct-1999 15:02:18 CDT, Signals from Hotfilm sheath down for awhile
16 Oct
The signals from the hotfilm sheath are down. They will be
down for a test of an analog breakout box.



85: Hotfilm, Site tower, Tue 12-Oct-1999 14:06:10 CDT,
9 Oct  1100 to 1700 Tried to find spiking problem.
9 Oct 1715  Steve/Gordon changed timing of zero circuit relative to sampling of sensors,  seems to take care of spiking problem. 
After this time the zero circuit is left on.
Before this time the zero circuit was off and base voltage was left to drift.
It was reset most days by re-installing circuit for short time. 
Spectra show spikes at 60 and 80 hz in the two cold films but not in the hot films. Spikes are small - we will probably remove them from the data after the fact.
10 Oct - 1145 to 1240 - reset timer on zero sheath covers on hotfilms.  Timer has been set wrong at zero 0300 and 1100 and 1200 hours.  Now set to cover sensors at hh:59 each hour.